# Finite-state transducers for substitution tilings

### v1

@@ 3 Relation to regular languages
Figure 5: Substitution system for Penrose P3 whole tiles
This proof fails in a substitution system with spurs, because in the case $\mathbf{I}_{i_{1},e}\mapsto\mathbf{E}_{u,v},\mathbf{I}_{i^{\prime}_{1},e^{\prime}}\mapsto\mathbf{E}_{u^{\prime},v^{\prime}}$, the edges $u,u^{\prime}$ of the two supertiles need not meet directly – they could be geometrically adjacent, but combinatorially separated by a spur. Worse, this combinatorial separation can translate into geometric separation in the next layer of supertiles: in a system with spurs, the supertiles of two adjacent tiles need not be adjacent, because they could be separated by some part of a third supertile which deflates into a spur to bring the original tiles together.
An extreme example of this occurs in the system shown in Figure 5, which describes the Penrose P3 tiling of thin and thick rhombs. This substitution system can be derived from a simpler one based on Robinson triangles in a similar manner to the way that the system in Figure 3 for the P2 tiling was derived from Figure 1.
Figure 6 shows an example of a succession of subtile/supertile relationships in that system. It begins on the left with a pair of thick rhomb tiles adjacent along one edge, and after passing from the tiles to their supertiles three times (or after deflating three times, if you read the diagram starting from the right-hand end), ends with another pair of thick rhombs adjacent along the same pair of edges – but, in between, passes through a state containing two tiles touching only at a vertex, and another state containing two tiles which do not touch at all.
A sequence of pairs of rhombs from successive deflations of a Penrose P3 tiling. In the first and last picture are two rhombs adjacent along an edge; the intermediate states include two rhombs touching only at a vertex, and two not touching at all. Between each adjacent pair of those diagrams, a diagram above shows how one position relates to the other, by showing the right-hand pair of rhombs each deflated into its subtiles, and highlighting the particular pair of subtiles that are shown on their own as the left-hand pair.
Figure 6: States in the P3 system with pairs of tiles not adjacent along an edge
If we were to try to build a DFA for this P3 substitution system, it would have to include at least the two states shown in Figure 6, and several more like them. But if the possible relationships between two $n$th-order supertiles are not restricted to the finite number of ways the two tiles can share an edge, then is there even a finite set of them at all?
In the case of this P3 system, there is a finite number of possible supertile relationships, and hence, the neighbour language $\mathcal{N}$ for this tiling is still regular. To show this, we can deflate each shape of Penrose rhomb twice into a patch of second-order subtiles with a more complex outline, and then imagine replacing the original tile shapes with the outlines of those patches. The twice-deflated tile shapes still tile the plane in the same way as the Penrose rhombs, but they extend beyond the boundaries of the original tile.
Using these twice-deflated shapes, it turns out that every possible edge-to-edge adjacency between tiles which can appear in a P3 tiling has the property that their $n$th-order supertiles are still adjacent along a whole edge, because in cases like Figure 6, the twice-deflated outlines extend beyond the original rhombus outlines and touch each other. Figure 7 shows the transformed version of each of the non-adjacent cases in Figure 6.
On the left, the thin and thick Penrose P3 rhombs are shown under double deflation, so that the thick rhomb turns into eight subtiles in total, and the thin one into five. On the right, two of the rhomb pairs from the previous figure -- the ones touching at a vertex and not touching at all -- are shown under the same transformation, demonstrating that in each case the doubly-deflated diagrams now include a shared edge between the deflations of the two rhombs.
Figure 7: Doubly-deflated P3 tile outlines (left) still touch in the states from Figure 6 (right)
Let us say that a substitution system preserves adjacency if, for any two tiles $t,t^{\prime}$ with distinct supertiles $s,s^{\prime}$, the tiles $t,t^{\prime}$ share an edge if and only if $s,s^{\prime}$ do, and the same is true for $n$th-order supertiles, for all $n$.
Then the DFA construction described in Theorem 1 can be extended to any substitution system which preserves adjacency. Spurless substitution systems always preserve adjacency; the P2 and P3 whole-tile substitution systems shown in Figures 3 and 5 do not, but each one can be turned into a system which does preserve adjacency by deflating their tile outlines twice.
All substitution systems studied by the author have a similar property: the ones described already for Penrose P2 and P3 tilings, various substitutions for the hat and Spectre tilings, the Ammann-Beenker tiling, and several simpler substitution systems for tile types that also admit periodic tilings. Therefore, it seems a natural conjecture that this always works:
Conjecture 1.
For any substitution system, there exists $n\geqslant 0$ such that deflating the tiles’ edges $n$ times makes it adjacency-preserving.
Conjecture 1 would imply that the neighbour language $\mathcal{N}$ is always regular. However, it is not necessary to prove it for all systems in order to work with a particular tiling, since it is easy to verify for each tiling one wants to work with.
What would happen if a substitution system did not satisfy this conjecture, so that no $n$-times deflation of the tile edges would make every possible relationship between two supertiles into an edge-adjacency? If the deflation of tile edges is iterated infinitely many times, then the tile shapes typically converge to a fractal, such as those shown in [Baake_2025] figure 3 for the hat HTPF system. Any relationship between two supertiles must make these fractals adjacent along some portion of their edge. So the only thing that can go wrong is that there are infinitely many ways for those fractals to be adjacent, all reachable as an intermediate state between two adjacent lowest-order tiles and some shared supertile. In this situation the neighbour language $\mathcal{N}$ would not be regular, because any attempt to recognise it with a DFA would omit any pair of adjacent tile addresses passing through a supertile relationship not represented by any of the DFA’s states.
@@
We show one more useful property of $\mathcal{N}$, related to reversibility. A DFA that recognises a regular language $L$ can always be turned into a recogniser for the language $L^{R}$ consisting of the reversal of every string of $L$, by simply reversing all the arrows in the state transition diagram and interchanging the starting and accepting states. Usually the reversed recogniser is no longer deterministic, and must be determinised again if a DFA for $L^{R}$ is desired. [10.1145/322326.322334] defines the concept of a zero-reversible acceptor to be a DFA for a language $L$ which remains deterministic under this reversal, because there is a unique accepting state, and no two distinct states have transitions to the same state on the same symbol. A zero-reversible language is one that admits a zero-reversible acceptor.
@@
In a substitution system that preserves adjacency, $\mathcal{N}(L,t)$ is zero-reversible.
@@
From the proof of Theorem 1, we can directly construct a DFA for $\mathcal{N}$ in which each non-accepting state describes two tiles $s,s^{\prime}$ in layer $L$ adjacent along edges $e,e^{\prime}$. (If the system is already adjacency-preserving, then $e,e^{\prime}$ describe edges of the actual tiles; if the tile outlines required one or more deflations to make it preserve adjacency, then $e,e^{\prime}$ describe edges of those deflated tile outlines. Either way, there are a finite number of ways two particular tile types can be adjacent, and each state describes one of them.) Accepting states describe a single tile $s$ in layer $L$. We denote these types of state as $S(L,s,e,s^{\prime},e^{\prime})$ and $A(L,s)$ respectively.
In some cases, the geometry of the tiles can force two tiles to be adjacent under multiple edges if they are adjacent along one. For example, tile $s$ might have edges $e,f$, and tile $s^{\prime}$ edges $e^{\prime},f^{\prime}$, such that the edges $e,e^{\prime}$ of the two tiles meet iff the edges $f,f^{\prime}$ meet. In this situation we consider $S(L,s,e,s^{\prime},e^{\prime})$ and $S(L,s,f,s^{\prime},f^{\prime})$ to be different names for the same DFA state.
@@
A transition to the accepting state is of the form $S(L,s,e,s^{\prime},e^{\prime})\to A(P,t)$, on the symbol $\Bigl(\bigl(\xleftarrow{i}t\bigr),\bigl(\xleftarrow{i^{\prime}}t\bigr)\Bigl)$. Such a transition is only valid if subtiles $i$ and $i^{\prime}$ of $t$ have types $s,s^{\prime}$ respectively. So all valid predecessors of $A(P,t)$ on this symbol have $L,s,s^{\prime}$ in common, and can only vary in the edge indices. But if those subtiles of $t$ are adjacent along multiple edges, say both $(e,e^{\prime})$ and $(f,f^{\prime})$, then we regarded $S(L,s,e,s^{\prime},e^{\prime})$ and $S(L,s,f,s^{\prime},f^{\prime})$ as the same state anyway. So $A(P,t)$ has at most one predecessor for any symbol.
Similarly, consider a transition between two non-accepting states $S(L,s,f,s^{\prime},f^{\prime})\to S(Q,u,e,u^{\prime},e^{\prime})$ on the symbol $\Bigl(\bigl(\xleftarrow{i}u\bigr),\bigl(\xleftarrow{i^{\prime}}u^{\prime}\bigr)\Bigr)$. Again, $s$ is constrained to be the tile type of subtile $i$ of $u$, and $s^{\prime}$ similarly. So all predecessors of $S(Q,u,e,u^{\prime},e^{\prime})$ on this symbol have $L,s,s^{\prime}$ in common and differ only in the edge indices, and again, by construction, all state descriptions referring to any valid $S(L,s,f,s^{\prime},f^{\prime})$ are names for the same state. ∎
@@ 4 Algorithms for handling addresses
In the previous section we showed (subject to Conjecture 1) that the combinatorics of a tiling can be handled by finite state automata. In this section we present concrete algorithms for constructing two different forms of automaton. The ultimate aim is to create a deterministic finite-state transducer which takes a tile edge address as input, and produces its neighbour edge address as output.
@@ 4.2 Computing neighbours using a transducer
In this section we shall assume that we are given a substitution system which is not a counterexample to Conjecture 1: it either preserves adjacency already, or can be made to do so by $n$-times deflation of all the tile edges, for some $n$.
@@ 4.2.1 Recogniser for neighbour edge pairs
One way to construct this DFA would be to follow the structure of the proof of Theorem 1: make a state for every possible adjacent pair of tiles, together with start and the required accepting states, and fill in the transitions. If the result is missing any transitions, transform the substitution system by deflating the edges of the tiles and try again.
However, deflating the tile edge shapes requires using the geometry of the tiling, which complicates implementation (the coordinates may well involve algebraic numbers, requiring either elaborate exact computation or risky floating-point approximation), as well as requiring more information as input. In this section we instead present a purely combinatorial algorithm to construct the same DFA from only the data described in section 2.5, based on Theorem 2 which tells us that $\mathcal{N}$ is the union of a set of zero-reversible languages.
In [10.1145/322326.322334] an algorithm is given to infer a DFA for a zero-reversible language from a finite corpus of strings of the language. Paraphrased, one first constructs an NFA consisting of a single start state, an accept state, and a separate path from start to accept for each string in the corpus. Then merge [^5] The description in [10.1145/322326.322334] suggests constructing the initial DFA in the form of a trie, so that states with the same predecessor start off already merged. This is equivalent. any two states $A,B$ which have transitions $U\xrightarrow{s}A,U\xrightarrow{s}B$ from the same predecessor $U$ on the same symbol $s$, and conversely, merge any two states with transitions to the same successor on the same symbol, i.e. $A\xrightarrow{s}V,B\xrightarrow{s}V$ for some $s,V$. Iterate both of those merging operations to closure. The first policy guarantees to make the NFA deterministic; the second makes its reversal deterministic too, so that the resulting DFA is still a DFA if the arrows are reversed and start swapped with accept.
@@
We have assumed in this section that the substitution system admits a regular neighbour language. Therefore, there is some finite length $k$ such that every state of a minimal DFA for that language is reached by some edge address pair of length $\leqslant k$. However, this $k$ is not known in advance, and may be greater than the length of strings we generated for our input corpus. So the final part of our algorithm is to detect if the output DFA is complete, and if not, retry with a corpus of longer address pairs. If $\mathcal{N}$ is regular at all, then sooner or later, an attempt will succeed.
@@ 4.2.2 Deterministic transducer
Algorithm 6 shows this procedure in pseudocode.
@@ 4.3.2 Neighbourhood-based refinement
A simple substitution system showing a ‘chair’, or L-triomino, subdivided into four smaller chairs. The sub-chairs are labelled L,R,B,C and the edges of the original chair are numbered from 0 to 7.
@@
A natural definition of the ‘neighbourhood’ of a tile in a tiling is to list, for each of the tile’s edges, what type of tile is adjacent to that edge, and which edge of the neighbour tile. We show an example of classifying tiles by this type of neighbourhood, based on the hierarchical substitution system of L-triominoes (or ‘chairs’) from [Goodman-Strauss1999] with a single tile type. Figure 8 shows the substitution rule for the chair tiling. It also numbers all the edges of the chair tile, and assigns identifiers for the four subtiles in a deflation ($B,L,R,C$ for ‘bottom’, ‘left’, ‘right’ and ‘centre’). Figure 9 shows the seven possible neighbourhoods of a single chair that can arise in a tiling generated by this system.
@@
To see why this helps, here is a concrete reason why the original chair tiling is ambiguous. Consider a tile which is the $B$ subtile of its supertile, which is the $B$ subtile of its supertile in turn, and so on $n$ times. That is, its address is of the form
$\textrm{chair}\underbrace{\xleftarrow{B}\textrm{chair}\xleftarrow{B}\textrm{chair}\xleftarrow{B}\cdots\xleftarrow{B}\textrm{chair}\xleftarrow{B}}_{\textstyle n\hbox{ steps}}\textrm{chair}$
so that the lowest-order tile occupies the very bottom corner of the $n$th-order supertile. What borders that tile on one of its lower edges – say, the one numbered 0 in Figure 8? It depends on what borders the $n$th-order supertile, and the various possibilities do not all lead to the same orientation of the smallest chair: this $n$th-order chair might be either back-to-back with the next one or nestled inside it, and the two touching lowest-order chairs will have the same relationship as the high-order ones in each case. So a transducer for this substitution system cannot generate even the first output symbol without looking at the $(n+1)$th input symbol. But $n$ was arbitrary. So the transducer cannot have finitely many states.
The refined chair tiling resolves this problem by requiring the input tile address to be specified in more detail. The deflation diagrams in Figure 11 show that the $B$ subtile of any chair has either type 3 or type 7: chair types 1,2,3 have a $B$ subtile of type 3, and types 4,5,6,7 have a $B$ subtile of type 7. In particular, each of types 3 and 7 has a $B$ subtile of the same type as itself. So the input address must now be specified in one of the following forms:
@@
However, a single run of the refinement algorithm is not always sufficient. The Ammann-Beenker tiling ([10.5555/19304], Figures 10.4.14 and 10.4.15) is an example of a substitution system requiring two passes of this refinement algorithm before a transducer can be successfully generated. Figure 12 shows the substitution system in a form suitable for use with these algorithms. The Ammann-Beenker tiling includes a constraint about the order of tiles that can appear around each vertex, shown in the figure as shaded markings which are required to form into a thick arrow shape around the vertex. The markings on the rhombus are asymmetric, and in order to meet the constraint, both mirror-image forms of it must occur in a specific pattern around any vertex where 8 rhombus tiles meet, shown in Figure 13.
@@
The more times this algorithm is run on a system, the more each tile type incorporates information from a larger and larger area of the tiling around the tile. It therefore seems reasonable to make the following conjecture:
Conjecture 2.
For any substitution system, there exists $n\geqslant 0$ such that applying Algorithm 9 $n$ times makes it unambiguous.

### v2

@@ 3 Relation to regular languages
Figure 5: Converging to fractal tile outlines for the P2 whole-tiles system
On the left, a group of six tiles is shown, consisting of the deflations of two larger kite tiles. The dart subtile of the top kite shares an edge with a kite subtile of the bottom kite, but the supertiles themselves only touch at a vertex. On the right, the same arrangement is shown with the fractal versions of the tile boundaries, and both the subtiles and supertiles share an identical segment of fractal boundary.
Figure 6: Adjacent tiles may not have adjacent supertiles, but the fractal versions do
In systems meeting the constraints described in section 2.1, iterating the geometric deflation process and re-scaling each deflated tile shape to the same overall area as the original tile causes the tile outlines to converge to a fractal. An example, based on the P2 whole-tile substitution system in Figure 3, is shown in Figure 5. Transforming the substitution system into this form, with every tile having a fractal outline and every tile edge being a fragment of that outline, restores the property that the deflation of a tile occupies the same plane region as the tile itself. Hence, when two tiles are adjacent along a segment of their outline, their two supertiles must also be: see Figure 6 for an example case. So we can still construct a set of automaton states based on two tiles sharing a segment of boundary.
We must also show that there are finitely many such states, which is no longer obvious when all we know is that two tiles share some segment of boundary instead of one of a discrete set of edges.
To see this, we observe that the fractal tile shapes are bounded. In a tiling of the plane using tiles from some particular layer, there are only finitely many ways to tile a disc of radius $r$ with a given tile $t$ at a fixed location in the centre, and the fractal forms of those patches exhibit all the ways that the fractal outline of tile $t$ can be adjacent to that of any other tile. No other forms of adjacency need be considered, because even if two fractal tile outlines $t,u$ physically fit together in some other way, that adjacency cannot occur between a pair of supertiles within the context of a legal tile address, since any such pair of supertiles form part of a tiling of the whole plane. ∎
Calculating the fractal tile shapes and their adjacencies is potentially awkward. We shall show a further useful property of $\mathcal{N}$, which we shall use in the next section to derive an easier, purely combinatorial algorithm for finding a recogniser for $\mathcal{N}$.
@@
The extra property relates to reversibility. A DFA that recognises a regular language $L$ can always be turned into a recogniser for the language $L^{R}$ consisting of the reversal of every string of $L$, by simply reversing all the arrows in the state transition diagram and interchanging the starting and accepting states. Usually the reversed recogniser is no longer deterministic, and must be determinised again if a DFA for $L^{R}$ is desired. [2] defines the concept of a zero-reversible acceptor to be a DFA for a language $L$ which remains deterministic under this reversal [^6] Why ‘zero-reversible’ and not just ‘reversible’? [2] defines the weaker condition of $k$-reversibility to mean that the reversed automaton need not be deterministic already, but if multiple transitions on the same symbol are available then $k$ extra symbols of lookahead suffice to rule out all but one of them. ‘Reversible’ by itself is used to mean ‘$k$-reversible for some $k\geqslant 0$’., because there is a unique accepting state, and no two distinct states have transitions to the same state on the same symbol. A zero-reversible language is one that admits a zero-reversible acceptor.
@@
$\mathcal{N}(L,t)$ is zero-reversible.
@@
From the proof of Theorem 1, we can directly construct a DFA for $\mathcal{N}$ in which each non-accepting state describes two tiles $s,s^{\prime}$ in layer $L$ adjacent along subsets $b,b^{\prime}$ of their boundary. (In a system with spurs, these boundary subsets may not correspond to single edges of the original tiles, but nonetheless, there will be a finite set of possible boundary subsets that $s,s^{\prime}$ can legally share.) Accepting states describe a single tile $s$ in layer $L$. We denote these types of state as $S(L,s,b,s^{\prime},b^{\prime})$ and $A(L,s)$ respectively.
We consider $b,b^{\prime}$ to be the maximal subsets of each tile’s boundary shared with the other’s. In some substitution systems, the geometry of two tiles means that if they are adjacent along some pair of edges $e,e^{\prime}$ then they are also adjacent along another edge pair $f,f^{\prime}$. For these purposes we regard $b,b^{\prime}$ as being the union of both of those edges, and any further edges the two tiles might share. Thus, we do not consider $S(L,s,e,s^{\prime},e^{\prime})$ and $S(L,s,f,s^{\prime},f^{\prime})$ to be distinct DFA states.
@@
A transition to the accepting state is of the form $S(L,s,b,s^{\prime},b^{\prime})\to A(P,t)$, on the symbol $\Bigl(\bigl(\xleftarrow{i}t\bigr),\bigl(\xleftarrow{i^{\prime}}t\bigr)\Bigl)$. Such a transition is only valid if subtiles $i$ and $i^{\prime}$ of $t$ have types $s,s^{\prime}$ respectively. Then subtiles $i,i^{\prime}$ of $t$ have a fixed geometric relationship, and there is only one maximal pair of boundary segments $b,b^{\prime}$ along which they can meet. So $A(P,t)$ has at most one predecessor for any symbol.
Similarly, consider a transition between two non-accepting states $S(M,s,b,s^{\prime},b^{\prime})\to S(Q,t,c,t^{\prime},c^{\prime})$ (where $Q$ is the parent layer of $M$) on the symbol $\Bigl(\bigl(\xleftarrow{i}t\bigr),\bigl(\xleftarrow{i^{\prime}}t^{\prime}\bigr)\Bigr)$. Again, $s$ is constrained to be the tile type of subtile $i$ of $u$, and $s^{\prime}$ similarly; and placing the fractal tile shapes of $t,t^{\prime}$ next to each other along boundary segments $c,c^{\prime}$ fixes their relative position in the plane, and leaves only one possibility for which segments $b,b^{\prime}$ of those subtiles can meet. ∎
@@ 4 Algorithms for handling addresses
In the previous section we showed that the combinatorics of a tiling can be handled by finite state automata. In this section we present concrete algorithms for constructing two different forms of automaton. The ultimate aim is to create a deterministic finite-state transducer which takes a tile edge address as input, and produces its neighbour edge address as output.
@@ 4.2.1 Recogniser for neighbour edge pairs
One way to construct this DFA would be to follow the structure of the proof of Theorem 1: make a state for every possible adjacent pair of tiles or their fractalisations, together with start and the required accepting states, and fill in the transitions by considering the geometry of the tiles.
However, using the geometry of the tiling complicates implementation (the coordinates may well involve algebraic numbers, requiring either elaborate exact computation or risky floating-point approximation), as well as requiring more information as input. In this section we instead present a purely combinatorial algorithm to construct the same DFA from only the data described in section 2.6, based on Theorem 2 which tells us that $\mathcal{N}$ is the union of a set of zero-reversible languages.
In [2] an algorithm is given to infer a DFA for a zero-reversible language from a finite corpus of strings of the language. Paraphrased, one first constructs an NFA consisting of a single start state, an accept state, and a separate path from start to accept for each string in the corpus. Then merge [^7] The description in [2] suggests constructing the initial DFA in the form of a trie, so that states with the same predecessor start off already merged. This is equivalent. any two states $A,B$ which have transitions $U\xrightarrow{s}A,U\xrightarrow{s}B$ from the same predecessor $U$ on the same symbol $s$, and conversely, merge any two states with transitions to the same successor on the same symbol, i.e. $A\xrightarrow{s}V,B\xrightarrow{s}V$ for some $s,V$. Iterate both of those merging operations to closure. The first policy guarantees to make the NFA deterministic; the second makes its reversal deterministic too, so that the resulting DFA is still a DFA if the arrows are reversed and start swapped with accept.
@@
By Theorem 1, the whole neighbour language $\mathcal{N}$ is regular. Therefore, there is some finite length $k$ such that every state of a minimal DFA for $\mathcal{N}$ is reached by some edge address pair of length $\leqslant k$. However, this $k$ is not known in advance, and may be greater than the length of strings we generated for our input corpus. So the final part of our algorithm is to detect if the output DFA is complete, and if not, retry with a corpus of longer address pairs. Since $\mathcal{N}$ is regular, sooner or later, an attempt will succeed.
@@
A full example of a recogniser generated by this algorithm is shown in Figure 7. In this case, since the substitution system is simple enough that adjacent tiles always have adjacent supertiles, it is possible to label each state of the recogniser with a specific pair of tile edges. In more complicated recognisers this is not always possible.
@@ 4.2.2 Deterministic transducer
Algorithm 6 shows this procedure in pseudocode. An example transducer built by the algorithm is shown in Figure 8.
A finite state machine derived from the Penrose P2 tiling, in its Robinson-triangle form. There is a single start state, four accepting states labelled with the triangle types, and 18 intermediate states labelled with pairs of triangle types and edge indices. Each transition is labelled with a pair of edge address symbols.
Figure 7: Example recogniser for neighbour edge pairs, in the P2 Robinson triangle system from Figure 22.
Non-accepting states are labelled with two tile types and the edges of those tiles that meet.
Accepting states (double outline) are labelled with a single tile type.
A second finite state machine derived from the Penrose P2 Robinson-triangle tiling. There is a single start state, four accepting states labelled with the triangle types, and 26 intermediate states. Unlike the previous state machine, each transition is labelled with one input edge address symbol and a string of output symbols with lengths ranging from 0 to 3. Each intermediate state is labelled with a set of states of the previous machine, with each of those states accompanied by another string of unemitted output symbols.
Figure 8: Example transducer for the P2 Robinson triangle system (Figure 22).
States other than start and accepting states are annotated with a set of (recogniser state, pending output) pairs, one on each line. The recogniser states are the same as the ones shown in Figure 7.
@@ 4.3.2 Neighbourhood-based refinement
A simple substitution system showing a ‘chair’, or L-triomino, subdivided into four smaller chairs. The sub-chairs are labelled 0,1,2,3 and the edges of the original chair are numbered from 0 to 7.
@@
A natural definition of the ‘neighbourhood’ of a tile in a tiling is to list, for each of the tile’s edges, what type of tile is adjacent to that edge, and which edge of the neighbour tile. We show an example of classifying tiles by this type of neighbourhood, based on the hierarchical substitution system of L-triominoes (or ‘chairs’) from [7] with a single tile type. Figure 9 shows the substitution rule for the chair tiling. It also numbers all the edges of the chair tile, and assigns identifiers for the four subtiles in a deflation ($B,L,R,C$ for ‘bottom’, ‘left’, ‘right’ and ‘centre’). Figure 10 shows the seven possible neighbourhoods of a single chair that can arise in a tiling generated by this system.
@@
To see why this helps, here is a concrete reason why the original chair tiling is ambiguous. Consider a tile which is subtile 0 of its supertile, which is subtile 0 of its supertile in turn, and so on $n$ times. That is, its address is of the form
$\textrm{chair}\underbrace{\xleftarrow{0}\textrm{chair}\xleftarrow{0}\textrm{chair}\xleftarrow{0}\cdots\xleftarrow{0}\textrm{chair}\xleftarrow{0}}_{\textstyle n\hbox{ steps}}\textrm{chair}$
so that the lowest-order tile occupies the very bottom corner of the $n$th-order supertile, in the orientation shown in figure 9. What borders that tile on one of its lower edges – say, edge 0 from the same figure? It depends on what borders the $n$th-order supertile, and the various possibilities do not all lead to the same orientation of the smallest chair: this $n$th-order chair might be either back-to-back with the next one or nestled inside it, and the two touching lowest-order chairs will have the same relationship as the high-order ones in each case. So a transducer for this substitution system cannot generate even the first output symbol without looking at the $(n+1)$th input symbol. But $n$ was arbitrary. So the transducer cannot have finitely many states.
The refined chair tiling resolves this problem by requiring the input tile address to be specified in more detail. The deflation diagrams in Figure 12 show that subtile 0 of any chair has either type 3 or type 7: for chair types 1,2,3, subtile 0 has type 3, and for types 4,5,6,7, it has type 7. In particular, each of types 3 and 7 has a subtile 0 of the same type as itself. So the input address must now be specified in one of the following forms:
@@
However, a single run of the refinement algorithm is not always sufficient. The Ammann-Beenker tiling ([8], Figures 10.4.14 and 10.4.15) is an example of a substitution system requiring two passes of this refinement algorithm before a transducer can be successfully generated. Figure 13 shows the substitution system in a form suitable for use with these algorithms. The Ammann-Beenker tiling includes a constraint about the order of tiles that can appear around each vertex, shown in the figure as shaded markings which are required to form into a thick arrow shape around the vertex. The markings on the rhombus are asymmetric, and in order to meet the constraint, both mirror-image forms of it must occur in a specific pattern around any vertex where 8 rhombus tiles meet, shown in Figure 14.
@@
The more times this algorithm is run on a system, the more each tile type incorporates information from a larger and larger area of the tiling around the tile. After some finite number of refinements, each tile will describe a neighbourhood of itself sufficiently large that, given any full tile address containing those refined tile types, the regions of the plane forced by the neighbourhoods of larger and larger supertiles expand in all directions and cover the whole plane. So it should always be possible to make a substitution system unambiguous after applying the refinement algorithm enough times.